OpenStudy (anonymous):

I'm having trouble verifying this identity. - tan^2x + sec^2x = 1

OpenStudy (anonymous):

tan^2 + 1 = sec^2x

OpenStudy (anonymous):

?

OpenStudy (anonymous):

wouldn't it be -tan^2(x) + sec^2(x) = 1?

OpenStudy (anonymous):

so if you subtract tan^2x from both sides what do you get?

OpenStudy (anonymous):

We know that tan^2x + 1 = sec^2x

OpenStudy (anonymous):

but that isn't verifying if i subtract sec^2(x)?

OpenStudy (anonymous):

-tan^2(x) + sec^2(x) -(sin(x)/cos(x))^2 + 1/cos^2(x) 1/cos^2(x) (-sin^2(x) + 1)) 1/cos^2(x)(cos^2(x)) = 1 is this correct?

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{- tan^2x + sec^2x = 1\\ -\dfrac{sin^2x}{cos^2x}+\dfrac{1}{cos^2x}\\~\\\ =\dfrac{-sin^2x+1}{cos^2x}\\~\\ =\dfrac{cos^2x}{cos^2x}\\ =1 }\end{align}\)